### v2

# Differentially-Private Distributed Model Predictive Control of Linear Discrete-Time Systems with Global Constraints

Abstract
Distributed model predictive control (DMPC) has attracted extensive attention as it can explicitly handle system constraints and achieve optimal control in a decentralized manner. However, the deployment of DMPC strategies generally requires the sharing of sensitive data among subsystems, which may violate the privacy of participating systems. In this paper, we propose a differentially-private DMPC algorithm for linear discrete-time systems subject to coupled global constraints. Specifically, we first show that a conventional distributed dual gradient algorithm can be used to address the considered DMPC problem but cannot provide strong privacy preservation. Then, to protect privacy against the eavesdropper, we incorporate a differential-privacy noise injection mechanism into the DMPC framework and prove that the resulting distributed optimization algorithm can ensure both provable convergence to a global optimal solution and rigorous $\epsilon$-differential privacy. In addition, an implementation strategy of the DMPC is designed such that the recursive feasibility and stability of the closed-loop system are guaranteed. Simulation results are provided to demonstrate the effectiveness of the developed approach.
Index Terms:
Distributed model predictive control, privacy preservation, differential privacy.
I Introduction
Over the past decades, model predictive control (MPC) has enjoyed great success due to its ability to explicitly handle system constraints and guarantee prescribed control performance [1, 2]. MPC can be implemented either in a centralized or distributed manner. Centralized MPC relies on a central unit to process all system information and solve the online optimization problem, which often results in poor scalability and requires substantial computation power, especially for complex or large-scale systems. Consequently, distributed MPC (DMPC) has garnered much attention in recent years, offering many advantages of distributed systems and proving to be an effective tool for various applications, including vehicle platoons [3], microgrids [4], and multi-robot systems [5].
Based on the type of couplings between subsystems, DMPC studies can be roughly divided into three categories, i.e., coupling in cost functions, coupling in system dynamics, and coupling in constraints. In this paper, we focus on systems with coupled global constraints, which have many real-world applications [6, 7]. Several approaches have been proposed to guarantee the strict satisfaction of coupled constraints in a distributed manner. In [8], a sequential DMPC method is developed, which first divides the global problem into several local subproblems and then solves them sequentially in a given order. The satisfaction of coupled constraints is guaranteed through plan exchanges among subsystems. [9] proposes a DMPC approach for flexible communication that can remove optimization order restrictions. Building on [8] and [9], [10] presents a parallel DMPC approach allowing simultaneous optimizations and maintaining flexible communication and parallel computation advantages. However, global optimality remains unclear in these approaches. In [11], a dual decomposition-based DMPC approach is designed to explicitly pursue global optimality. This approach transforms the dual problem of DMPC into a consensus optimization problem and then solves it using the distributed alternating direction multiplier method (ADMM). To improve the convergence speed when solving DMPC, a Nesterov-accelerated-gradient algorithm is utilized in [12]. In addition, a push-sum dual gradient algorithm and a noisy ADMM algorithm are developed in [13] and [14] to address the DMPC problem under time-varying directed communication network and communication noise, respectively.
In the aforementioned methods [11, 12, 13, 14], distributed optimization algorithms are leveraged to address the DMPC problem, requiring each subsystem/agent to share explicit local information with neighboring subsystems/agents in order to satisfy coupled global constraints. Note that the shared messages often contain sensitive information, which raises significant concerns about privacy leakage. For instance, an eavesdropper could wiretap the communication channel and deduce privacy-sensitive information from exchanged messages. When DMPC is adopted in specific domains like smart grid or intelligent transportation, the disclosure of privacy-sensitive information may further pose safety risks and lead to economic losses. Considering the growing awareness of privacy and security, it is imperative to ensure privacy protection in DMPC. So far, few results are available on the privacy preservation of DMPC, but privacy-preserving approaches for distributed optimization have been well studied. For the latter, one typical technique is (partially) homomorphic encryption, which has been utilized in existing works such as [15, 16]. The encryption-based approaches use cryptography to conceal privacy-sensitive information and can be directly extended for the privacy protection of DMPC [17]. However, this technique generally requires tedious encryption and decryption procedures, leading to huge overheads in both communication and computation. Another technique relies on spatially or temporally correlated noises/uncertainties [18, 19, 20], aiming to obscure information shared in distributed optimization. Due to the correlated nature of these noises/uncertainties, these approaches are typically vulnerable to adversaries with access to all messages shared in the communication network.
With implementation simplicity and rigorous mathematical foundations, differential privacy (DP) has witnessed growing popularity, emerging as a de facto standard for privacy protection. In recent years, DP-based privacy methods have been introduced in distributed optimization by integrating DP noise into objective functions [21] or exchanged information [22, 23, 24]. Nevertheless, the direct injection of persistent DP noise to existing algorithms inevitably compromises optimization performance, resulting in an inherent trade-off between accuracy and privacy. It is crucial to note that extending DP-based privacy approaches to DMPC is not straightforward, as the compromise on optimization accuracy can deteriorate control performance and potentially lead to constraint violations.
In this paper, a differentially-private DMPC algorithm is designed for linear discrete-time systems with coupled global constraints. We first demonstrate the need for privacy preservation by showcasing that a conventional distributed dual gradient algorithm for DMPC is vulnerable to eavesdropping attacks. A DP noise injection mechanism is then introduced into the distributed dual gradient algorithm, which obscures the private information exchanged among subsystems to prevent adversaries from inferring sensitive information. By leveraging the results in [25, 26], a weakening factor sequence and a step-size sequence are carefully designed to effectively mitigate the influence of DP noise. Rigorous analysis shows that the proposed algorithm can ensure almost sure convergence to a global optimal solution and maintain $\epsilon$-differential privacy with a finite cumulative privacy budget. Aligned with the privacy-preserving distributed algorithm, we provide an implementation strategy for DMPC, ensuring the recursive feasibility and stability of the closed-loop system. Simulations are performed to validate the performance of the proposed scheme.
The rest of this paper is organized as follows. In Section II, the preliminaries of DMPC and differential privacy are introduced. In Section III, a new differentially-private distributed dual gradient algorithm is developed, and convergence analysis is conducted. Section IV presents the implementation strategy of DMPC. Finally, a numerical study is given in Section V, and concluding remarks are summarized in Section VI.
Notations: $\mathbb{R}^{n}$ stands for the $n$-dimensional Euclidean space. Given two integers $a$ and $b$ ($a<b$), $\mathbb{Z}_{a}^{b}$ represents the set $\{a,a+1,\cdots,b\}$. $I_{n}$ denotes the identify matrix of dimension $n$. ${\bf 1}_{n}$ and ${\bf 0}_{n}$ represent the $n$-dimensional column vector with all entries being 1 and 0, respectively. We use $Q>(\geq)0$ to denote that $Q$ is a positive definite (semi-definite) matrix. $\|x\|$ and $\|x\|_{1}$ represent the standard Euclidean norm and the $L_{1}$ norm of a vector $x$, respectively. Moreover, $\|x\|_{Q}^{2}:=x^{\top}Qx$.
II Problem Formulation and Preliminaries
II-A Problem Description
Consider $M$ linear discrete-time subsystems where each is described as follows:
$x_{i}(t+1)=A_{i}x_{i}(t)+B_{i}u_{i}(t),\quad i\in\mathbb{Z}_{1}^{M}.$ (1)
In (1), $x_{i}(t)\in\mathbb{R}^{n_{i}}$ and $u_{i}(t)\in\mathbb{R}^{m_{i}}$ are the state and control input of subsystem $i$ at time instant $t$, respectively. The state and control input of subsystem $i$ should satisfy the following local constraints:
$x_{i}(t)\in\mathcal{X}_{i},\quad u_{i}(t)\in\mathcal{U}_{i},$ (2)
where $\mathcal{X}_{i}\subset\mathbb{R}^{n_{i}}$ and $\mathcal{U}_{i}\subset\mathbb{R}^{m_{i}}$ denote state and control input constraint sets, respectively. Moreover, all the subsystems are subject to $p$ global constraints described by
$\sum_{i=1}^{M}(\varPsi_{x_{i}}x_{i}(t)+\varPsi_{u_{i}}u_{i}(t))\leq{\bf 1}_{p},$ (3)
where $\varPsi_{x_{i}}\in\mathbb{R}^{p\times n_{i}}$ and $\varPsi_{u_{i}}\in\mathbb{R}^{p\times m_{i}}$ are some given matrices.
Assumption 1.
Each linear discrete-time subsystem, i.e., $(A_{i},B_{i})$, is controllable. Additionally, $\mathcal{X}_{i}$ and $\mathcal{U}_{i}$ are bounded and closed polytopes which contain the origins as their inner point.
In this paper, we consider the same DMPC problem as presented in [11, 12, 13, 14, 17]. Specifically, based on (1)-(3), the DMPC problem is formulated as
$\mathcal{P}:\quad$ $\min_{\{\tilde{{\bm{u}}}_{1},\cdots,\tilde{{\bm{u}}}_{M}\}}\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})$ (4a)
$\;\mathrm{s.t.}\quad\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$ (4b)
$\quad\quad\sum_{i=1}^{M}f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})\leq b(\varepsilon).$ (4c)
In (4a), $J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})$ is the local objective function, which is defined as
$J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})\!:=\!\sum_{\ell=0}^{N-1}\!\left(\|\tilde{x}_{i}(\ell|t)\|_{Q_{i}}^{2}\!+\!\|\tilde{u}_{i}(\ell|t)\|_{R_{i}}^{2}\right)\!+\!\|\tilde{x}_{i}(N|t)\|_{P_{i}}^{2},$ (5)
where $N\in\mathbb{Z}_{>0}$ is the length of prediction horizon, $\tilde{x}_{i}(\ell|t)$ and $\tilde{u}_{i}(\ell|t)$ are the $\ell$th step predicted state and control input at time instant $t$, respectively, $\tilde{{\bm{u}}}_{i}:=\{\tilde{u}_{i}(0|t),\cdots,\tilde{u}_{i}(N-1|t)\}$ stands for the predicted input sequence over the prediction horizon, and $Q_{i}>0$, $R_{i}>0$, and $P_{i}>0$ are weight matrices. For each subsystem $i$, $P_{i}$ is the solution of the following algebraic Riccati equation:
$(A_{i}+B_{i}K_{i})^{\top}P_{i}(A_{i}+B_{i}K_{i})-P_{i}=-(Q_{i}+K_{i}^{\top}R_{i}K_{i}),$ (6)
where $K_{i}:=-(R_{i}+B_{i}^{\top}P_{i}B_{i})^{-1}B_{i}^{\top}P_{i}A_{i}$. The local constraint set $\tilde{\mathcal{U}_{i}}(x_{i}(t))$ in (4b) is formulated as
$\tilde{\mathcal{U}_{i}}(x_{i}(t)):=\{\tilde{{\bm{u}}}_{i}\in\mathbb{R}^{m_{i}N}:$ (7)
$\tilde{x}_{i}(\ell+1|t)=A_{i}\tilde{x}_{i}(\ell|t)+B_{i}\tilde{u}_{i}(\ell|t),\tilde{x}_{i}(0|t)=x_{i}(t),$
$\tilde{x}_{i}(\ell|t)\in\mathcal{X}_{i},\tilde{u}_{i}(\ell|t)\in\mathcal{U}_{i},\tilde{x}_{i}(N|t)\in\mathcal{X}_{i}^{f},\ell\in\mathbb{Z}_{0}^{N-1}\},$
with $\mathcal{X}_{i}^{f}$ being the terminal constraint set. In addition, the global coupled constraint in (4c) is a tightened form of the constraint in (3), and $f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})$ and $b(\varepsilon)$ are given by
$f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})$ $:=\begin{bmatrix}\varPsi_{x_{i}}\tilde{x}_{i}(0|t)+\varPsi_{u_{i}}\tilde{u}_{i}(0|t)\\ \vdots\\ \varPsi_{x_{i}}\tilde{x}_{i}(N-1|t)+\varPsi_{u_{i}}\tilde{u}_{i}(N-1|t)\end{bmatrix},$ (8)
$b(\varepsilon)$ $:=\begin{bmatrix}(1-\varepsilon M){\bf 1}_{p}\\ \vdots\\ (1-\varepsilon MN){\bf 1}_{p}\end{bmatrix},$
where $0\leq\varepsilon<\frac{1}{MN}$ is a tolerance parameter. The introduction of the tightened constraint is to ensure that the numerical algorithm used to solve the DMPC problem can be terminated in advance. To facilitate the feasibility and stability analysis of DMPC, the terminal constraint set $\mathcal{X}_{i}^{f}$ can be chosen as a closed maximal polytope such that for any $x_{i}\in\mathcal{X}_{i}^{f}$, we have
$K_{i}x_{i}\in\mathcal{U}_{i},\quad(A_{i}+B_{i}K_{i})x_{i}\in\mathcal{X}_{i}^{f},$ (9)
$\sum_{i=1}^{M}(\varPsi_{x_{i}}+\varPsi_{u_{i}}K_{i})x_{i}\leq(1-\varepsilon MN){\bf 1}_{p}.$
For more details about the tightening of constraint (3) and the construction of the terminal constraint set $\mathcal{X}_{i}^{f}$, please refer to [11].
Assumption 2.
For the initial system state $\{x_{1}(0),\cdots,x_{M}(0)\}$, the Slater condition holds, i.e., there exists $\{\tilde{{\bm{u}}}_{1},\cdots,\tilde{{\bm{u}}}_{M}\}$ that satisfies (4b) and (4c).
The communication network of $M$ subsystems is described by an interaction weight matrix $L=\{L_{ij}\}\in\mathbb{R}^{M\times M}$. Specifically, for each subsystem $i$, the neighbor set $\mathcal{N}_{i}$ consists of all subsystems $j$ that can directly communicate with subsystem $i$. If $j\in\mathcal{N}_{i}$, then $L_{ij}>0$; otherwise, $L_{ij}=0$. We define $L_{ii}:=-\sum_{j\in\mathcal{N}_{i}}L_{ij}$ for all $i\in\mathbb{Z}_{1}^{M}$. Moreover, $L$ satisfies the following assumption:
Assumption 3.
The interaction weight matrix $L=\{L_{ij}\}$ is symmetric and satisfies ${\bf 1}_{M}^{\top}L={\bf 0}_{M}^{\top}$, $L{\bf 1}_{M}={\bf 0}_{M}$, and $\|I_{M}+L-\frac{{\bf 1}_{M}{\bf 1}_{M}^{\top}}{M}\|<1$.
Assumption 3 guarantees that the communication network described by $L$ is connected, meaning that there exists a path from any subsystem to any other subsystem.
II-B Distributed Dual-Gradient Method
The Lagrangian function corresponding to the optimization problem in (4) is given by
($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\lambda)$ $\!=\!\sum_{i=1}^{M}\!J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})\!+\!\lambda^{\top}\left(\sum_{i=1}^{M}f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})-b(\varepsilon)\!\right)$ (10)
$=\sum_{i=1}^{M}\left(J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})+\lambda^{\top}g_{i}(\tilde{{\bm{u}}}_{i})\right),$
where $\lambda\in\mathbb{R}^{Np}_{+}$ (the non-negative orthant of $\mathbb{R}^{Np}$) is the Lagrangian multiplier and $g_{i}(\tilde{{\bm{u}}}_{i}):=f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})-\frac{b(\varepsilon)}{M}$. The dual problem of (4) is defined as
$\max_{\lambda\geq 0}\min_{\{\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))\}}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\lambda).$ (11)
Based on Assumptions 1, 2 and the definition of DMPC problem, it can be concluded that the strong duality holds for (4), and the optimization problem (4) can be addressed by solving its dual problem (11). In addition, the Saddle-Point Theorem holds, i.e., given an optimal primal-dual pair ($(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$, the following relationship holds for any $\lambda\in\mathbb{R}^{Np}_{+}$ and $\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$:
($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda)\leq\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\leq\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\lambda^{*}).$ (12)
A conventional approach to solving problem (11) is the distributed dual-gradient method [27, 28]. The core idea is to regard the Lagrangian multiplier (dual variable) $\lambda$ as a consensus variable and then $M$ subsystems address the optimization problem in a collective manner. Specifically, let each subsystem have a local copy $\lambda_{i}^{k}$ of the dual variable. $\Pi_{\mathbb{R}^{Np}_{+}}\left[\cdot\right]$ denotes Euclidean projection of a vector on the set +$\mathbb{R}^{Np}_{+}$, and $\gamma^{k}>0$ denotes the step-size. Then, the distributed dual-gradient method is summarized in Algorithm 1, and the overall DMPC implementation is detailed in Algorithm 2.
Input: $x_{i}(t)$, $i\in\mathbb{Z}_{1}^{M}$
Output: $\tilde{{\bm{u}}}_{i}^{\bar{k}}$, $i\in\mathbb{Z}_{1}^{M}$
1 Initialization: set $\lambda_{i}^{0}\in\mathbb{R}^{Np}_{+}$ and $\tilde{{\bm{u}}}_{i}^{0}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$, $\forall i\in\mathbb{Z}_{1}^{M}$ Parameters: deterministic sequence $\gamma^{k}>0$
2for $k=0,2,\cdots,\bar{k}-1$ do
3       for all $i\in\mathbb{Z}_{1}^{M}$ (in parallel) do
4             Every subsystem $i$ sends $\lambda_{i}^{k}$ to subsystem $j\in\mathcal{N}_{i}$;
5             After receiving $\lambda_{j}^{k}$ from all $j\in\mathcal{N}_{i}$, subsystem $i$ updates its primal and dual variables:
$\tilde{\lambda}_{i}^{k}=\lambda_{i}^{k}+\sum_{j\in\mathcal{N}_{i}}L_{ij}(\lambda_{j}^{k}-\lambda_{i}^{k});$ (13)
$\tilde{{\bm{u}}}_{i}^{k+1}=\mathop{\rm argmin}_{\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})+(\tilde{\lambda}_{i}^{k})^{\top}g_{i}(\tilde{{\bm{u}}}_{i});$ (14)
$\lambda_{i}^{k+1}\!=\Pi_{\mathbb{R}^{Np}_{+}}\left[\tilde{\lambda}_{i}^{k}+\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right];$ (15)
6       end for
7
8 end for
Algorithm 1 Distributed Dual-gradient Algorithm
1 At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
2Every subsystem $i$ computes $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ by following Algorithm 1 with $x_{i}(t)$;
3Set the input sequence as $\tilde{{\bm{u}}}_{i}(t)=\tilde{{\bm{u}}}_{i}^{\bar{k}}$;
4Apply $\tilde{u}_{i}(0|t)$ to subsystem $i$;
Wait for the next time instant; let $t=t+1$ and go to step 1.
Algorithm 2 DMPC Algorithm
If Assumptions 2 and 3 hold and if the step-size $\gamma^{k}$ satisfies the conditions $\sum_{k=0}^{\infty}\gamma^{k}=\infty$, $\sum_{k=0}^{\infty}(\gamma^{k})^{2}<\infty$, then Algorithm 1 guarantees the convergence of the sequence $\{\tilde{{\bm{u}}}_{i}^{k}\}$, i.e., $\lim_{k\rightarrow\infty}\|\tilde{{\bm{u}}}_{i}^{k}-\tilde{{\bm{u}}}_{i}^{*}\|$. Note that the objective function of problem (4) is strictly convex, and thus the asymptotic primal convergence can be established without resorting to local averaging mechanism (see Section 3.4 in [28] for more details). In addition, if the iteration number $\bar{k}$ is selected sufficiently large to meet specific criteria [29, 11], then Algorithm 2 ensures the feasibility and stability of the considered MPC problem.
In Algorithm 1, each subsystem can avoid sharing the primal variable and only share its local copy $\lambda_{i}^{k}$ of the dual variable with its neighbors. However, this sharing mechanism cannot provide strong privacy protection as the iteration trajectory of $\lambda_{i}^{k}$ still bears information of the primal variable. In particular, we assume that the adversary has prior knowledge about the communication network $L$ and the step-size $\gamma^{k}$, and can get access to all information exchanged in communication channels. Under this circumstance, the adversary can record the updates of $\bar{\lambda}_{i}^{k}$ and $\lambda_{i}^{k}$ at each iteration. Then, based on $\bar{\lambda}_{i}^{k}$ and $\lambda_{i}^{k+1}$ in two consecutive iterations and $\gamma^{k}$, the adversary can employ (15) to estimate the value of $g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})$. It should be noted that $g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})$ is privacy-sensitive as it is the function of primal variable and is used to formulate the coupled global constraint. Therefore, it is necessary to incorporate a privacy protection mechanism into the distributed dual-gradient algorithm such that the DMPC problem can be addressed with privacy protection.
II-C On Differential Privacy
In this work, DP is used to characterize and quantify the achieved privacy level of distributed optimization algorithms. Given the continual exchange of information among subsystems in iterative optimization algorithms, the notion of $\epsilon$-DP for continuous bit streams [30] is adopted. Drawing inspiration from the distributed optimization framework proposed by [22], we represent the DMPC problem in (4) by four parameters $(L,\mathcal{J},\tilde{\mathcal{U}},\mathcal{G})$ to facilitate DP analysis. Specifically, $L$ is the interaction weight matrix describing the communication network, $\mathcal{J}:=\{J_{1},\,\cdots,J_{M}\}$ denotes the set of objective functions for individual subsystems, $\tilde{\mathcal{U}}:=\{\tilde{\mathcal{U}_{1}},\cdots,\tilde{\mathcal{U}_{M}}\}$ is the domain of optimization variables, and $\mathcal{G}:=\{g_{1},\cdots,g_{M}\}$ represents the set of constraint functions for individual subsystems. The adjacency between two optimization problems is defined as follows:
Definition 1.
Two distributed optimization problems $\mathcal{P}=(L,\mathcal{J},\tilde{\mathcal{U}},\mathcal{G})$ and $\mathcal{P}^{{}^{\prime}}=(L^{{}^{\prime}},\mathcal{J}^{{}^{\prime}},\tilde{\mathcal{U}}^{{}^{\prime}},\mathcal{G}^{{}^{\prime}})$ are adjacent if they satisfy the following conditions:
  • the interaction weight matrices, the objective functions, and the domains of optimization variables are identical, i.e., $L=L^{{}^{\prime}}$, $\mathcal{J}=\mathcal{J}^{{}^{\prime}}$, and $\tilde{\mathcal{U}}=\tilde{\mathcal{U}}^{{}^{\prime}}$;
  • there exists an $i\in\mathbb{Z}_{1}^{M}$ such that $g_{i}\neq g_{i}^{{}^{\prime}}$, and $g_{j}=g_{j}^{{}^{\prime}}$ for all $j\in\mathbb{Z}_{1}^{M},\,j\neq i$;
  • $g_{i}$ and $g_{i}^{{}^{\prime}}$, while different, exhibit similar behaviors near $\theta^{*}$, where $\theta^{*}$ is the solution of $\mathcal{P}$. More precisely, there exists a $\delta>0$ such that for all $\bm{u}_{i}$ and $\bm{u}_{i}^{{}^{\prime}}$ within the domain $B_{\delta}(\theta^{*}):=\{\bm{v}:\bm{v}\in\mathbb{R}^{Nm_{i}},\|\bm{v}-\theta^{*}\|<\delta\}$, $g_{i}(\bm{u}_{i})=g_{i}^{{}^{\prime}}(\bm{u}_{i}^{{}^{\prime}})$ holds.
We denote the execution of a distributed optimization algorithm as $\mathcal{A}$, represented by a sequence of the iteration variable $\vartheta$, i.e., $\mathcal{A}=\{\vartheta^{0},\vartheta^{1},\cdots\}$. We assume that adversaries have access to all communicated messages among the subsystems. Hence, under an execution $\mathcal{A}$, the adversaries’ observation is the sequence of communicated messages, denoted as $\mathcal{O}$. Let $\mathbb{O}$ represent the set of all possible observation sequences. For a distributed optimization problem $\mathcal{P}$ with an initial state $\vartheta^{0}$, the observation mapping is defined as $\mathcal{R}_{\mathcal{P},\vartheta^{0}}(\mathcal{A}):=\mathcal{O}$. Furthermore, for a distributed optimization problem $\mathcal{P}$, an initial state $\vartheta^{0}$, and an observation sequence $\mathcal{O}$, $\mathcal{R}_{\mathcal{P},\vartheta^{0}}^{-1}(\mathcal{O})$ denotes the set of executions $\mathcal{A}$ capable of generating the observation $\mathcal{O}$.
Definition 2 ($\epsilon$-differential privacy, [22]).
For a given $\epsilon>0$, an iterative distributed algorithm ensures $\epsilon$-differential privacy if for any two adjacent optimization problems $\mathcal{P}$ and $\mathcal{P}^{\prime}$, any initial state ${\vartheta}^{0}$, and any set of observation sequences $\mathcal{O}_{s}\subseteq\mathbb{O}$, the following relationship always holds:
$\mathbb{P}[\mathcal{R}_{\mathcal{P},\vartheta^{0}}\left(\mathcal{O}_{s}\right)]\leq e^{\epsilon}\mathbb{P}[\mathcal{R}_{\mathcal{P}^{{}^{\prime}},\vartheta^{0}}\left(\mathcal{O}_{s}\right)],$ (16)
with the probability $\mathbb{P}$ taken over the randomness of iteration processes.
The definition of $\epsilon$-DP guarantees that adversaries, with access to all communicated information, cannot infer knowledge about any participating subsystem’s sensitive information. It can be found that a smaller $\epsilon$ indicates a better extent of privacy preservation.
III Differentially-Private Distributed Dual-Gradient Algorithm
III-A Algorithm Description
In this section, a DP noise injection mechanism is proposed to achieve privacy preservation in the distributed dual-gradient algorithm. The developed algorithm is summarized in Algorithm 3.
In contrast to Algorithm 1, where each subsystem directly sends $\lambda_{i}^{k}$ to its neighbors, Algorithm 3 incorporates DP noise $\zeta_{i}^{k}$ into $\lambda_{i}^{k}$ and shares the perturbed signal $\hat{\lambda}_{i}^{k}:=\lambda_{i}^{k}+\zeta_{i}^{k}$ among the communication network. Therefore, the information available to potential adversaries is the sequence $\{\hat{\lambda}_{i}^{k}\}$. Due to the randomness of DP noise, it is impossible for the adversary to extract useful information from $\{\hat{\lambda}_{i}^{k}\}$ with significant probability. Furthermore, it should be noted that directly integrating persistent DP noise into existing optimization algorithms will compromise the convergence accuracy. To address this issue, we utilize findings from [25, 26] to design a weakening factor. As shown in (18), the weakening factor, denoted as $\chi^{k}$, is applied on the interaction terms ($L_{ij}(\hat{\lambda}_{j}^{k}-\lambda_{i}^{k})$). The fundamental principle behind incorporating this weakening factor is to gradually eliminate the impact of DP noise on convergence accuracy.
Input: $x_{i}(t)$, $i\in\mathbb{Z}_{1}^{M}$
Output: $\tilde{{\bm{u}}}_{i}^{\bar{k}}$, $i\in\mathbb{Z}_{1}^{M}$
1 Initialization: set $\lambda_{i}^{0}\in\mathbb{R}^{Np}_{+}$ and $\tilde{{\bm{u}}}_{i}^{0}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$, $\forall i\in\mathbb{Z}_{1}^{M}$ Parameters: deterministic sequence $\gamma^{k}>0$ and $\chi^{k}>0$
2for $k=0,2,\cdots,\bar{k}-1$ do
3       for all $i\in\mathbb{Z}_{1}^{M}$ (in parallel) do
4             Every subsystem $i$ adds DP noise $\zeta_{i}^{k}$ to $\lambda_{i}^{k}$, and then sends the obscured value $\hat{\lambda}_{i}^{k}:=\lambda_{i}^{k}+\zeta_{i}^{k}$ to subsystem $j\in\mathcal{N}_{i}$;
5             After receiving $\hat{\lambda}_{j}^{k}$ from all $j\in\mathcal{N}_{i}$, subsystem $i$ updates its primal and dual variables:
$\tilde{{\bm{u}}}_{i}^{k+1}=\mathop{\rm argmin}_{\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})+(\lambda_{i}^{k})^{\top}g_{i}(\tilde{{\bm{u}}}_{i});$ (17)
[…]
6       end for
7
8 end for
Algorithm 3 Differentially-private Distributed Dual-gradient Algorithm
To facilitate the convergence and privacy analysis, the following DP noise assumption is introduced:
Assumption 4.
For every $k$ and every $i\in\mathbb{Z}_{1}^{M}$, conditional on $\lambda_{i}^{k}$, the DP noise $\zeta_{i}^{k}$ satisfies $\mathbb{E}\left[\zeta_{i}^{k}\mid\lambda_{i}^{k}\right]=0$ and $\mathbb{E}\left[\|\zeta_{i}^{k}\|^{2}\mid\lambda_{i}^{k}\right]=(\sigma_{i}^{k})^{2}$ for all $k\geq 0$, and
$\sum_{k=0}^{\infty}(\chi^{k})^{2}\,\max_{i\in\mathbb{Z}_{1}^{M}}(\sigma_{i}^{k})^{2}<\infty,$ (19)
where $\{\chi^{k}\}$ is the weakening factor sequence from Algorithm 3.
Considering Assumption 4, we use the Laplace noise mechanism to generate $\zeta_{i}^{k}$ and then add it to all shared messages. More specifically, given a constant $\nu>0$, let ${\rm Lap}(\nu)$ represent a Laplace distribution of a scalar random variable, and $\rho\rightarrow\frac{1}{2\nu}e^{-\frac{|\rho|}{\nu}}$ be the corresponding probability density function. At each iteration $k$, every element of $\zeta_{i}^{k}$ is independently sampled from Laplace distribution ${\rm Lap}(\nu^{k})$, where $\nu^{k}>0$. One can verify that the mean and variance of ${\rm Lap}(\nu^{k})$ is zero and $2(\nu^{k})^{2}$, respectively. Therefore, $\zeta_{i}^{k}$ satisfies $\mathbb{E}\left[\zeta_{i}^{k}\mid\lambda_{i}^{k}\right]=0$ and $\mathbb{E}\left[\|\zeta_{i}^{k}\|^{2}\mid\lambda_{i}^{k}\right]=(\sigma_{i}^{k})^{2}=2(\nu^{k})^{2}$.
Remark 1.
In Algorithm 3, we allow the variance of DP noise $\zeta_{i}^{k}$, i.e., $2(\nu^{k})^{2}$, to be constant or increasing with $k$. To satisfy condition (19), one can carefully design the weakening factor sequence $\{\chi^{k}\}$ to make its decreasing rate outweigh the increasing rate of the noise level sequence $\{\nu^{k}\}$. For instant, (19) can be satisfied by setting $\chi^{k}=\frac{c_{1}}{1+c_{2}k^{c_{3}}}$ and $\nu^{k}=d_{1}+d_{2}k^{d_{3}}$ with any $c_{1}>0$, $c_{2}>0$, $10.5<c_{3}<$ 1, $d_{1}>0$, $d_{2}>0$, and $0<d_{3}<0.5-c_{3}$.
III-B Convergence Analysis
The arithmetic average of local dual variables $\lambda_{i}^{k}$ is given by
$\bar{\lambda}^{k}=\frac{1}{M}\sum_{i=1}^{M}\lambda_{i}^{k}.$ (20)
The relation between $\lambda_{i}^{k}$ and $\bar{\lambda}^{k}$ is summarized in the following theorem.
Theorem 1.
Suppose Assumptions 1, 3, and 4 hold. If the non-negative weakening factor sequence $\{\chi^{k}\}$ and the step-size sequence $\{\gamma^{k}\}$ in Algorithm 3 satisfy
$\sum_{k=0}^{\infty}\chi^{k}=\infty,\,\sum_{k=0}^{\infty}(\chi^{k})^{2}<\infty,\,\sum_{k=0}^{\infty}\frac{(\gamma^{k})^{2}}{\chi^{k}}<\infty,$ (21)
then the following results hold almost surely:
  1. 1.
    $\lim_{k\to\infty}\left\|\lambda_{i}^{k}-\bar{\lambda}^{k}\right\|=0$ for all $i\in\mathbb{Z}_{1}^{M}$;
  2. 2.
    $\sum_{k=0}^{\infty}\chi^{k}\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\bar{\lambda}^{k}\right\|^{2}<\infty$;
  3. 3.
    $\sum_{k=0}^{\infty}\gamma^{k}\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\bar{\lambda}^{k}\right\|<\infty$.
Proof.
Based on Assumption 1 and (7), it can be concluded that the local constraint set $\tilde{\mathcal{U}_{i}}(x_{i}(t))$ is bounded. Then, from (8) and the relation $g_{i}(\tilde{{\bm{u}}}_{i}):=f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})-\frac{b(\varepsilon)}{M}$, we have that for any $\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$, $g_{i}(\tilde{{\bm{u}}}_{i})$ is bounded, i.e., there exists a constant $C_{g}\in\mathbb{R}_{+}$ such that
$\|g_{i}(\tilde{{\bm{u}}}_{i})\|\leq C_{g},\forall\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t)),i\in\mathbb{Z}_{1}^{M}.$ (22)
According to Assumptions 3, 4, (21), and (22), we can follow the same line of reasoning as that of Theorem 1 in [26] to obtain the results. ∎
The following lemma is also required for convergence analysis:
Lemma 1 (Lemma 11, [31]).
Let $\{\psi^{k}\}$, $\{\phi^{k}\}$, $\{a^{k}\}$, and $\{b^{k}\}$ be random non-negative scalar sequences such that
$\mathbb{E}\left[\psi^{k+1}|\mathcal{F}^{k}\right]\leq(1+{a^{k}})\psi^{k}-\phi^{k}+{b^{k}},\quad\forall k\geq 0,$
[…]
Theorem 2.
Suppose Assumptions 1, 3, and 4 hold. If the non-negative sequences $\{\chi^{k}\}$ and $\{\gamma^{k}\}$ satisfy $\sum_{k=0}^{\infty}\chi^{k}=\infty$, $\sum_{k=0}^{\infty}(\chi^{k})^{2}<\infty$, $\sum_{k=0}^{\infty}\gamma^{k}=\infty$, and $\sum_{k=0}^{\infty}\frac{(\gamma^{k})^{2}}{\chi^{k}}<\infty$, then Algorithm 3 guarantees that
$\lim_{k\rightarrow\infty}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})=\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*}),$
$\lim_{k\rightarrow\infty}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k}\},\lambda^{*})=\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$
hold almost surely.
Proof.
Based on Lemma 1 in [32] and the update law of $\lambda_{i}^{k}$ in (18), it can be obtained that for any $\lambda\in\mathbb{R}_{+}^{Np}$,
$\sum_{i=1}^{M}\left\|\lambda_{i}^{k+1}-\lambda\right\|^{2}$ (23)
$\leq$ $\sum_{i=1}^{M}\left\|\lambda_{i}^{k}+\chi^{k}\sum_{j\in\mathcal{N}_{i}}L_{ij}(\hat{\lambda}_{j}^{k}-\lambda_{i}^{k})+\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})-\lambda\right\|^{2}$
$\leq$ $\sum_{i=1}^{M}\left\|\lambda_{i}^{k}\!+\!\chi^{k}\sum_{j\in\mathcal{N}_{i}}L_{ij}(\lambda_{j}^{k}+\zeta_{j}^{k}-\lambda_{i}^{k})\!+\!\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})-\lambda\right\|^{2}$
$\leq$ $\sum_{i=1}^{M}\left\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\lambda+\chi^{k}\xi_{i}^{k}+\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right\|^{2}$
$\leq$ $\sum_{i=1}^{M}\left(\left\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\lambda\right\|^{2}+\left\|\chi^{k}\xi_{i}^{k}+\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right\|^{2}\right.$
$\qquad\left.+2\left(\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\lambda\right)^{\top}\left(\chi^{k}\xi_{i}^{k}\right)\right.$
$\qquad\left.+2\left(\bar{\lambda}^{k}-\lambda\right)^{\top}\left(\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right)\right.$
$\qquad\left.+2\left(\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\bar{\lambda}^{k}\right)^{\top}\left(\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right)\right),$
where $w_{ij}^{k}$ and $\xi_{i}^{k}$ are defined as
$w_{ii}^{k}:=1+\chi^{k}L_{ii},\quad w_{ij}^{k}:=\chi^{k}L_{ij},\quad\xi_{i}^{k}:=\sum_{j\in\mathcal{N}_{i}}L_{ij}\zeta_{j}^{k}.$ (24)
According to Assumptions 3, 4 and (24), one can verify that
$w_{ij}^{k}=w_{ji}^{k},\quad\sum_{i=1}^{M}w_{ij}^{k}=\sum_{j=1}^{M}w_{ij}^{k}=1,$ (25)
$\mathbb{E}\left[\xi_{i}^{k}\mid\lambda_{i}^{k}\right]=0,\quad\mathbb{E}\left[\|\xi_{i}^{k}\|^{2}\mid\lambda_{i}^{k}\right]=\sum_{j\in\mathcal{N}_{i}}(L_{ij}\sigma_{j}^{k})^{2}.$ (26)
Using (25), it can be derived that
$\sum_{i=1}^{M}\left\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\lambda\right\|^{2}$ $=\sum_{i=1}^{M}\left\|\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\left(\lambda_{j}^{k}-\lambda\right)\right\|^{2}$ (27)
$\leq\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\lambda\right\|^{2}.$
It can be obtained from (17) that for any $\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$, $J_{i}(x_{i}(t),\tilde{{\bm{u}}}^{k+1}_{i})+(\lambda_{i}^{k})^{\top}g_{i}(\tilde{{\bm{u}}}^{k+1}_{i})\leq J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})+(\lambda_{i}^{k})^{\top}g_{i}(\tilde{{\bm{u}}}_{i})$. Thus, we can further derive that
$\sum_{i=1}^{M}\left(\bar{\lambda}^{k}-\lambda\right)^{\top}\left(\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right)$ (28)
$=$ $\gamma^{k}\sum_{i=1}^{M}\left(\left(\bar{\lambda}^{k}-\lambda_{i}^{k}\right)^{\top}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})+\left(\lambda_{i}^{k}-\lambda\right)^{\top}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right.$
$\qquad\left.+J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{k+1})-J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{k+1})\right)$
$\leq$ $\gamma^{k}\sum_{i=1}^{M}\left(\left(\bar{\lambda}^{k}-\lambda_{i}^{k}\right)^{\top}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})+\left(\lambda_{i}^{k}-\bar{\lambda}^{k}\right)^{\top}g_{i}(\tilde{{\bm{u}}}_{i})\right.$
$\qquad\left.+J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})+\left(\bar{\lambda}^{k}\right)^{\top}g_{i}(\tilde{{\bm{u}}}_{i})\right.$
$\qquad\left.-J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{k+1})-\lambda^{\top}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right)$
$\leq$ $\gamma^{k}\sum_{i=1}^{M}\left(\left(\bar{\lambda}^{k}-\lambda_{i}^{k}\right)^{\top}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})+\left(\lambda_{i}^{k}-\bar{\lambda}^{k}\right)^{\top}g_{i}(\tilde{{\bm{u}}}_{i})\right)$
$+\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda)\right).$
Using (26)-(28) and the fact that $\|g_{i}(\tilde{{\bm{u}}}_{i})\|\leq C_{g}$, $\forall\tilde{{\bm{u}}}_{i}\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$, we can take the conditional expectation with respect to $\mathcal{F}^{k}=\{\lambda_{\ell}^{k},\tilde{\bm{u}}_{\ell}^{k+1};0\leq\ell\leq k\}$ in (23) to obtain
$\sum_{i=1}^{M}\mathbb{E}\left[\left\|\lambda_{i}^{k+1}-\lambda\right\|^{2}|\mathcal{F}^{k}\right]$ (29)
$\leq$ $\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\lambda\right\|^{2}\!+d^{k}\!+2\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda)\right),$
where $d^{k}$ is given by
$d^{k}=$ $(\chi^{k})^{2}\sum_{i=1}^{M}\sum_{j\in\mathcal{N}_{i}}(L_{ij}\sigma_{j}^{k})^{2}+M(\gamma^{k})^{2}C_{g}^{2}$ (30)
$+6C_{g}\gamma^{k}\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\bar{\lambda}^{k}\right\|.$
Based on Assumption 4, Theorem 1, and the conditions for $\chi^{k}$ and $\gamma^{k}$ in (21), it can be concluded that $d^{k}$ is summable, i.e., $\sum_{k=0}^{\infty}d^{k}<\infty$.
Plugging the optimal primal-dual pair ($(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$ into (29) and utilizing the Saddle-Point Theorem (12), we can arrive at
$\sum_{i=1}^{M}\mathbb{E}\left[\left\|\lambda_{i}^{k+1}-\lambda^{*}\right\|^{2}|\mathcal{F}^{k}\right]$ (31)
$\leq$ $\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\lambda^{*}\right\|^{2}\!+\!d^{k}\!+\!2\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\right),$
and
$\sum_{i=1}^{M}\mathbb{E}\left[\left\|\lambda_{i}^{k+1}-\lambda^{*}\right\|^{2}|\mathcal{F}^{k}\right]$ (32)
$\leq$ $\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\lambda^{*}\right\|^{2}\!+\!d^{k}\!+\!2\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\!-\!\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})\right).$
According to Lemma 1, (31), and (32), it can be concluded that the following relationships hold almost surely:
$\sum_{i=1}^{M}\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})\right)<\infty,$ (33)
$\sum_{i=1}^{M}\gamma^{k}\left(\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})\right)<\infty.$
Since $\gamma^{k}$ is non-summable, we have that ($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$ and ($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})$ converge to zero almost surely. ∎
Remark 2.
The conditions for the weakening factor sequence $\{\chi^{k}\}$ and the step-size sequence $\{\gamma^{k}\}$ in Theorems 1 and 2 can be satisfied, e.g., by selecting $\chi^{k}=\frac{c_{1}}{1+c_{2}k^{c_{3}}}$ and $\gamma^{k}=\frac{c_{4}}{1+c_{5}k}$ with any $c_{1}>0$, $c_{2}>0$, $10.5<c_{3}<$ 1, $c_{4}>0$, and $c_{5}>0$. Note that the design of $\chi^{k}$ in this example is identical to the one in Remark 1. Therefore, the sequences $\{\chi^{k}\}$, $\{\gamma^{k}\}$, and $\{\nu^{k}\}$ can be meticulously tailored to meet all conditions required by Assumption 4 and Theorems 1, 2.
III-C Privacy Analysis
Based on the adjacency concept delineated in Definition 1, we can establish two adjacent distributed optimization problems, denoted as $\mathcal{P}$ and $\mathcal{P}^{\prime}$. There is only one signal that differs between these two problems, and without loss of generality we denote it as $g_{i}$ in $\mathcal{P}$ and $g^{\prime}_{i}$ in $\mathcal{P}^{\prime}$. According to the third condition of Definition 1, signals $g_{i}$ and $g^{{}^{\prime}}_{i}$ are required to exhibit similar behaviors around the optimal solution, i.e., $g_{i}$ and $g^{{}^{\prime}}_{i}$ should converge to each other if the algorithm can ensure convergence to the optimal solution. Therefore, according to the proven convergence in Theorem 2, we can formalize this condition by stipulating the existence of a constant $C>0$ such that
$\|g_{i}(\tilde{\bm{u}}^{k}_{i})-{g^{{}^{\prime}}_{i}(\tilde{\bm{u}}^{\prime k}_{i})}\|_{1}\leq C\chi^{k}$ (34)
holds for all $k\geq 0$.
For Algorithm 3, an execution is represented as $\mathcal{A}=\{\vartheta^{0},\vartheta^{1},\ldots\}$ with $\vartheta^{k}=\lambda^{k}=\begin{bmatrix}(\lambda_{1}^{k})^{\top},\cdots,(\lambda_{M}^{k})^{\top}\end{bmatrix}^{\top}$. An observation sequence is denoted as $\mathcal{O}=\{o^{0},o^{1},\ldots\}$ with $o^{k}=\hat{\lambda}^{k}=\begin{bmatrix}(\hat{\lambda}_{1}^{k})^{\top},\cdots,(\hat{\lambda}_{M}^{k})^{\top}\end{bmatrix}^{\top}$ (note that $\hat{\lambda}_{i}^{k}=\lambda_{i}^{k}+\zeta_{i}^{k}$, as detailed in Algorithm 3). Similar to the sensitivity metric proposed for constraint-free distributed optimization in [22], we formulate the sensitivity of Algorithm 3 in the following manner:
Definition 3.
At each iteration $k$, for any two adjacent distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$ and any initial state $\vartheta^{0}$, the sensitivity of Algorithm 3 is given by
$\Delta^{k}:=\sup\limits_{\mathcal{O}\in\mathbb{O}}\left\{\sup\limits_{\vartheta\in\mathcal{R}_{\mathcal{P},\vartheta^{0}}^{-1}(\mathcal{O}),\>\vartheta^{{}^{\prime}}\in\mathcal{R}_{\mathcal{P}^{{}^{\prime}},\vartheta^{0}}^{-1}(\mathcal{O})}\|\vartheta^{k}-\vartheta^{\prime k}\|_{1}\right\},$ (35)
where $\mathbb{O}$ denotes the set of all possible observation sequences.
Given Definition 3, we have the following lemma:
Lemma 2.
In Algorithm 3, at each iteration $k$, if each subsystem’s DP noise vector $\zeta_{i}^{k}\in\mathbb{R}^{Np}$ comprises $Np$ independent Laplace noises with parameter $\nu^{k}$, satisfying $\sum_{k=1}^{T_{0}}\frac{\Delta^{k}}{\nu^{k}}\leq\bar{\epsilon}$ for some $\bar{\epsilon}>0$, then Algorithm 3 achieves $\epsilon$-differential privacy with the cumulative privacy level for iterations $0\leq k\leq T_{0}$ less than $\bar{\epsilon}$.
Proof.
The proof of this lemma follows the same reasoning as that of Lemma 2 in [22]. ∎
We also introduce the following lemma for privacy analysis:
Lemma 3.
(Lemma 4, [33]) Let $\{\psi^{k}\}$ be a non-negative sequence, and $\{a^{k}\}$ and $\{b^{k}\}$ be positive sequences satisfying $\sum_{k=0}^{\infty}a^{k}=\infty$, $\lim_{k\rightarrow\infty}a^{k}=0$, and $\frac{b^{k}}{a^{k}}$ converges to zero with a polynomial rate. If there exists a $\bar{K}\geq 0$ such that $\psi^{k+1}\leq(1-a^{k})\psi^{k}+b^{k}$ holds for all $k\geq\bar{K}$, then it follows that $\psi^{k}\leq\bar{C}\frac{b^{k}}{a^{k}}$ for all $k$, with $\bar{C}$ being some constant.
Theorem 3.
Suppose the conditions of Theorem 1 hold. If every element of $\zeta_{i}^{k}$ is independently sampled from Laplace distribution ${\rm Lap}(\nu^{k})$, where $(\sigma_{i}^{k})^{2}=2(\nu^{k})^{2}$ satisfies Assumption 4, then the following results hold:
  1. 1.
    For any finite number of iterations $T$, Algorithm 3 ensures $\epsilon$-differential privacy, and the cumulative privacy budget is bounded by $\epsilon\leq\sum_{k=1}^{T}\frac{C\varsigma^{k}}{\nu^{k}}$, where $\varsigma^{k}:=\sum_{s=1}^{k-1}\Pi_{q=s}^{k-1}(1-\chi^{q}\bar{L})\gamma^{s-1}\chi^{s-1}+\gamma^{k-1}\chi^{k-1}$, $\bar{L}:=\min_{i\in\mathbb{Z}_{1}^{M}}|L_{ii}|$, and $C$ is from (34);
  2. 2.
    If $\sum_{k=0}^{\infty}\frac{\gamma^{k}}{\nu^{k}}<\infty$ holds, then the cumulative privacy budget remains finite as $T\rightarrow\infty$.
Proof.
To establish the privacy guarantees, we begin by analyzing the sensitivity of Algorithm 3. Given any initial state $\lambda^{0}$, any fixed observation $\mathcal{O}$, and two adjacent distributed optimization problems $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$, the sensitivity depends on $\|\lambda^{k}-\lambda^{\prime k}\|_{1}$ as per Definition 3. Note that $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$ differ solely in one signal, and without loss of generality, we denote this distinct signal as the $i$th one, i.e., $g_{i}$ in $\mathcal{P}$ and $g^{{}^{\prime}}_{i}$ in $\mathcal{P}^{{}^{\prime}}$. Since the initial conditions and observations of $\mathcal{P}$ and $\mathcal{P}^{{}^{\prime}}$ are the same for $j\neq i$, it follows that $\lambda_{j}^{k}=\lambda^{\prime k}_{j}$ for all $k$ and $j\neq i$. Consequently, $\|\lambda^{k}-\lambda^{\prime k}\|_{1}$ is always equal to $\|\lambda_{i}^{k}-\lambda^{\prime k}_{i}\|_{1}$.
Based on (18) in Algorithm 3, $L_{ii}:=-\sum_{j\in\mathbb{N}_{i}}L_{ij}$, and the fact that the observations $\lambda_{j}^{k}+\zeta_{j}^{k}$ and ${\lambda^{\prime}_{j}}^{k}+{\zeta^{\prime}_{j}}^{k}$ are identical, we can derive that
$\|\lambda_{i}^{k+1}-{\lambda^{\prime}_{i}}^{k+1}\|_{1}\leq$ $(1-|L_{ii}|\chi^{k})\|\lambda_{i}^{k}-{\lambda^{\prime}_{i}}^{k}\|_{1}$ (36)
$\quad+\gamma^{k}\|g_{i}(\tilde{\bm{u}}_{i}^{k+1})-{g^{{}^{\prime}}_{i}(\tilde{\bm{u}}^{\prime k+1}_{i})}\|_{1}.$
Therefore, it can be obtained from (34) and (36) that the sensitivity $\Delta^{k}$ is bounded by
$\Delta^{k+1}$ $\leq(1-|L_{ii}|\chi^{k})\Delta^{k}+\gamma^{k}\|g_{i}(\tilde{\bm{u}}_{i}^{k+1})-{g^{{}^{\prime}}_{i}(\tilde{\bm{u}}^{\prime k+1}_{i})}\|_{1}$ (37)
$\leq(1-|L_{ii}|\chi^{k})\Delta^{k}+C\gamma^{k}\chi^{k}.$
According to Lemma 2 and (37), the first statement can be obtained.
Lemma 3 is exploited to prove the second statement of Theorem 3. Specifically, based on (37) and the properties of $\chi^{k}$ and $\gamma^{k}$, Lemma 3 can be used to conclude that there exists some constant $\bar{C}$ such that the sensitivity $\Delta^{k}$ satisfies $\Delta^{k}\leq\bar{C}\gamma^{k}$. It can be further obtained from Lemma 2 that $\epsilon\leq\sum_{k=1}^{T}\frac{\bar{C}\gamma^{k}}{\nu^{k}}$. Thus, if $\sum_{k=0}^{\infty}\frac{\gamma^{k}}{\nu^{k}}<\infty$ holds (i.e., the sequence $\{\frac{\gamma^{k}}{\nu^{k}}\}$ is summable), then $\epsilon$ will be finite even when $T\rightarrow\infty$. ∎
IV Implementation of Privacy-Preserving DMPC
In this section, the overall DMPC implementation is described based on the differentially-private distributed dual-gradient algorithm.
IV-A Algorithm Implementation
Algorithm 3 will terminate after $\bar{k}$ iterations. Note that Algorithm 3 converges almost surely in a probability sense, and thus the global constraints (3) may not be satisfied within a given number of iterations. Based on (8), one can verify that the global constraints are satisfied if the following condition holds:
$\sum_{i=1}^{M}g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})=\sum_{i=1}^{M}f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{\bar{k}})-b(\varepsilon)\leq\varepsilon M{\bf 1}_{Np}.$ (38)
To verify whether the global constraints are satisfied after the termination of Algorithm 3, we employ a privacy-preserving static average consensus method developed in [34].
Specifically, after Algorithm 3 terminates, each subsystem initializes $z_{i}^{0}=g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})=f_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{\bar{k}})-\frac{b(\varepsilon)}{M}$. Then, $z_{i}^{0}$ is decomposed into two substates $z_{i,\alpha}^{0}$ and $z_{i,\beta}^{0}$, where $z_{i,\alpha}^{0}$ and $z_{i,\beta}^{0}$ are randomly chosen from the set of all real numbers with the constraint $z_{i,\alpha}^{0}+z_{i,\beta}^{0}=2z_{i}^{0}$. The static average consensus method updates $z_{i,\alpha}^{\ell}$ and $z_{i,\beta}^{\ell}$ as follows:
$z_{i,\alpha}^{\ell+1}$ $=z_{i,\alpha}^{\ell}+\iota\sum_{j\in\mathcal{N}_{i}}a_{ij}^{\ell}(z_{j,\alpha}^{\ell}-z_{i,\alpha}^{\ell})+\iota a_{i,\alpha\beta}^{\ell}(z_{i,\beta}^{\ell}-z_{i,\alpha}^{\ell}),$ (39)
$z_{i,\beta}^{\ell+1}$ $=z_{i,\beta}^{\ell}+\iota a_{i,\alpha\beta}^{\ell}(z_{i,\alpha}^{\ell}-z_{i,\beta}^{\ell}),$
where $\iota$, $a_{i,\alpha\beta}^{\ell}$, $a_{i,j}^{\ell}\in\mathbb{R}_{+}$. As proven in [34], by appropriately selecting the parameters $\iota$, $a_{i,\alpha\beta}^{\ell}$, and $a_{i,j}^{\ell}$, $z_{i,\alpha}^{\ell}$ and $z_{i,\beta}^{\ell}$ converge to the average consensus value $\frac{1}{M}\sum_{i=1}^{M}z_{i}^{0}$ (i.e., $\frac{1}{M}\sum_{i=1}^{M}g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$). Therefore, each subsystem can utilize the converged value of $z_{i,\alpha}^{\ell}$ to check whether condition (38) is satisfied. It is worth noting that conventional static average consensus approaches [35, 36, 37] can also be employed to calculate the value of $\frac{1}{M}\sum_{i=1}^{M}z_{i}^{0}$ in a distributed manner. However, these approaches necessitate subsystems to directly share $z_{i}^{0}$ with their neighbors, potentially leading to privacy breaches as $z_{i}^{0}=g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$ contains sensitive information about $\tilde{{\bm{u}}}_{i}^{\bar{k}}$. The average consensus method developed in [34] employs a state decomposition scheme to mask the real values of $z_{i}^{0}$. As shown in (39), the substate $z_{i,\alpha}^{\ell}$ governs the role of internode interactions and is the only value from subsystem $i$ that can be seen by its neighbors. On the other hand, the other substate $z_{i,\beta}^{\ell}$ participates in the distributed interactions by solely interacting with $z_{i,\alpha}^{\ell}$. Hence, the existence of $z_{i,\beta}^{\ell}$ is invisible to neighboring nodes of subsystem $i$, although it directly affects the evolution of $z_{i,\alpha}^{\ell}$. Through this state decomposition design, strong privacy preservation can be guaranteed. For further details, please refer to [34].
After executing the static average consensus method, an update mechanism is designed for the control input sequence $\tilde{{\bm{u}}}_{i}(t)$. Based on the consensus results, if condition (38) is met, then the solution $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ at the current time instant is applied to $\tilde{{\bm{u}}}_{i}(t)$; otherwise, the control input sequence from the last time instant, i.e., $\tilde{{\bm{u}}}_{i}(t-1)$, is used to update $\tilde{{\bm{u}}}_{i}(t)$. The overall DMPC strategy is presented in Algorithm 4. By utilizing the static average consensus method and the update mechanism designed for $\tilde{{\bm{u}}}_{i}(t)$, we can ensure that if the solution $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ at time instant $t=0$ is feasible, then Algorithm 4 can generate feasible solutions for the remaining time.
1 At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
2Every subsystem $i$ computes $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ by following Algorithm 3 with $x_{i}(t)$;
3Every subsystem $i$ runs the static average consensus algorithm (39) to obtain $\sum_{i=1}^{M}g_{i}(\tilde{{\bm{u}}}_{i}^{\bar{k}})$;
4if Condition (38) is satisfied then
5      Set current control input sequence $\tilde{{\bm{u}}}_{i}(t):=\{\tilde{u}_{i}(0|t),\tilde{u}_{i}(1|t),\cdots,\tilde{u}_{i}(N-1|t)\}$ as $\tilde{{\bm{u}}}_{i}(t)=\tilde{{\bm{u}}}_{i}^{\bar{k}}$;
6else
7      Use $\tilde{{\bm{u}}}_{i}(t-1)$ to update $\tilde{{\bm{u}}}_{i}(t)$, i.e.,
$\tilde{{\bm{u}}}_{i}(t)=$$\{\tilde{u}_{i}(1|t-1),\tilde{u}_{i}(2|t-1),\cdots,$
$\quad\tilde{u}_{i}(N-1|t-1),K_{i}\tilde{x}_{i}(N|t-1)\};$
8 end if
9
10Save $\tilde{{\bm{u}}}_{i}(t)$ in subsystem $i$; apply $\tilde{u}_{i}(0|t)$ to subsystem $i$;
Wait for the next time instant; let $t=t+1$ and go to step 1.
Algorithm 4 Privacy-preserving DMPC Algorithm
IV-B Feasibility and Stability
At any time instant $t$, the solution $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ generated from Algorithm 3 is constrained in the bounded set $\tilde{\mathcal{U}_{i}}(x_{i}(t))$, and thus $J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i})$ is bounded and the following relation holds:
$\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{\bar{k}})-\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{*})\leq\eta,$ (40)
where $\eta\in\mathbb{R}_{+}$ is a bounded constant. The following theorem summarizes the theoretical results of the developed DMPC strategy.
Theorem 4.
Assume that $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ generated from Algorithm 3 satisfies the global constraints at time instant $t=0$. Then, the following results hold:
  1. 1.
    If Algorithm 4 has a feasible solution at time instant $t$, then it has a feasible solution at $t+1$.
  2. 2.
    If $\{x_{i}\in\mathbb{R}^{n_{i}}:\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}\subset\mathcal{X}^{f}_{i}$, then the state trajectory of each subsystem converges to the terminal set $\mathcal{X}^{f}_{i}$ in finite time.
Proof.
As shown in Algorithm 4, the input sequence at time instant $t$ is denoted by $\tilde{{\bm{u}}}_{i}(t)=\{\tilde{u}_{i}(0|t),\tilde{u}_{i}(1|t),\cdots,\tilde{u}_{i}(N-1|t)\}$. Let $\tilde{{\bm{x}}}_{i}(t)=\{\tilde{x}_{i}(0|t),\tilde{x}_{i}(1|t),\cdots,\tilde{x}_{i}(N|t)\}$ be the corresponding predicted state sequence. Since $\tilde{{\bm{u}}}_{i}(t)$ is a feasible solution, it can be obtained from (7), (8), and (38) that $\tilde{{\bm{u}}}_{i}(t)\in\tilde{\mathcal{U}_{i}}(x_{i}(t))$ and
$\sum_{i=1}^{M}\varPsi_{x_{i}}\tilde{x}_{i}(\ell|t)+\varPsi_{u_{i}}\tilde{u}_{i}(\ell|t)\leq(1-\varepsilon M\ell){\bf 1}_{p},\ell\in\mathbb{Z}_{0}^{N-1}.$ (41)
At time instant $t+1$, an input sequence $\hat{{\bm{u}}}_{i}(t+1)$ and its corresponding predicted state sequence $\hat{{\bm{x}}}_{i}(t+1)$ are defined as
$\hat{{\bm{u}}}_{i}(t+1)$ (42)
$=$ $\{\hat{u}_{i}(0|t+1),\hat{u}_{i}(1|t+1),\cdots,\hat{u}_{i}(N-1|t+1)\}$
$=$ $\{\tilde{u}_{i}(1|t),\tilde{u}_{i}(2|t),\cdots,\tilde{u}_{i}(N-1|t),K_{i}\tilde{x}_{i}(N|t)\},$
$\hat{{\bm{x}}}_{i}(t+1)$
$=$ $\{\hat{x}_{i}(0|t+1),\hat{x}_{i}(1|t+1),\cdots,\hat{x}_{i}(N|t+1)\}$
$=$ $\{\tilde{x}_{i}(1|t),\tilde{x}_{i}(2|t),\cdots,\tilde{x}_{i}(N|t),(A_{i}+B_{i}K_{i})\tilde{x}_{i}(N|t)\}.$
Based on (9), (41), and (42), it can be concluded that $\hat{{\bm{u}}}_{i}(t+1)\in\tilde{\mathcal{U}_{i}}(x_{i}(t+1))$ and
$\sum_{i=1}^{M}\varPsi_{x_{i}}\hat{x}_{i}(\ell|t+1)+\varPsi_{u_{i}}\hat{u}_{i}(\ell|t+1)$ (43)
$=$ $\sum_{i=1}^{M}\varPsi_{x_{i}}\tilde{x}_{i}(\ell+1|t)+\varPsi_{u_{i}}\tilde{u}_{i}(\ell+1|t)$
$\leq$ $(1-\varepsilon M(\ell+1)){\bf 1}_{p},\quad\ell\in\mathbb{Z}_{0}^{N-2},$
$\sum_{i=1}^{M}\varPsi_{x_{i}}\hat{x}_{i}(N-1|t+1)+\varPsi_{u_{i}}\hat{u}_{i}(N-1|t+1)$
$=$ $\sum_{i=1}^{M}(\varPsi_{x_{i}}+\varPsi_{u_{i}}K_{i})\tilde{x}_{i}(N|t)\leq(1-\varepsilon MN){\bf 1}_{p}.$
Therefore, $\hat{{\bm{u}}}_{i}(t+1)$ is a feasible solution at time instant $t+1$, which completes the proof for the first statement of Theorem 4. From the above analysis, it is evident that at $t=0$, if $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ generated from Algorithm 3 is feasible, then the update mechanism designed for $\tilde{{\bm{u}}}_{i}(t)$ in Algorithm 4 ensures the solution feasibility for the remaining duration.
To prove the second statement, we first define a Lyapunov function $V(\{x_{i}(t)\}):=\sum_{i=1}^{M}J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}^{*})$. According to the algebraic Riccati equation (6) and (42), we have
$J_{i}(x_{i}(t+1),\hat{{\bm{u}}}_{i}(t+1))-J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))$ (44)
$=$ $-\|x_{i}(t)\|^{2}_{Q_{i}}-\|\tilde{u}_{i}(0|t)\|^{2}_{R_{i}}.$
$\hat{{\bm{u}}}_{i}(t+1)$ is a feasible solution at $t+1$ but may not be optimal. Thus, we have
$V(\{x_{i}(t+1)\})$ (45)
$\leq$ $\sum_{i=1}^{M}J_{i}(x_{i}(t+1),\hat{{\bm{u}}}_{i}(t+1))$
$=$ $\sum_{i=1}^{M}\left(J_{i}(x_{i}(t),\tilde{{\bm{u}}}_{i}(t))-\|x_{i}(t)\|^{2}_{Q_{i}}-\|\tilde{u}_{i}(0|t)\|^{2}_{R_{i}}\right)$
$\leq$ $V(\{x_{i}(t)\})+\eta-\sum_{i=1}^{M}\|x_{i}(t)\|^{2}_{Q_{i}},$
where the equality condition is due to (44) and the last inequality follows from (40). (45) indicates that $x_{i}(t)$ converges to the bounded set $\{\{x_{i}\}:\sum_{i=1}^{M}\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}$ in finite time. Considering this fact and the assumption that $\{x_{i}\in\mathbb{R}^{n_{i}}:\|x_{i}\|^{2}_{Q_{i}}\leq\eta\}\subset\mathcal{X}^{f}_{i}$, it can be concluded that $x_{i}(t)$ enters the terminal set $x_{i}(t)$ in finite time. ∎
V Numerical Simulations
In this section, simulation is conducted to demonstrate the performance of the developed method. A group of four linear time-invariant subsystems are considered. The network structure of these four subsystems is shown in Figure 1. The system matrices $A_{i}$ and $B_{i}$ are chosen as
$A_{i}$ $=\begin{bmatrix}1&1\\ 0&1\end{bmatrix},B_{i}=\begin{bmatrix}1\\ 1\end{bmatrix},i=1,3,$
$A_{i}$ $=\begin{bmatrix}2&1\\ 0&1\end{bmatrix},B_{i}=\begin{bmatrix}1\\ 1\end{bmatrix},i=2,4.$
For all subsystems, the local state and input constraint sets are selected as $\mathcal{X}_{i}=\{x_{i}:-1\leq x_{i}\leq 1\}$ and $\mathcal{U}_{i}=\{u_{i}:-0.3\leq u_{i}\leq 0.3\}$, respectively. The global constraint is $-0.65\leq\sum_{i=1}^{4}u_{i}\leq 0.65$. The weight matrices $Q_{i}$ and $R_{i}$ are set as $Q_{i}=I$ and $R_{i}=0.1$, respectively. The length of the prediction horizon is chosen as $N=5$. In Algorithm 3, we inject Laplace noise with parameter $\nu^{k}=0.1+0.001k^{0.1}$. The weakening factor sequence and step-size sequence is set as $\chi^{k}=\frac{2}{1+0.01k^{0.9}}$ and $\gamma^{k}=\frac{5}{1+0.1k}$, respectively. In the simulation, Algorithm 4 is executed 20 times, and the mean and the variance of the state and input trajectories are computed. For comparison, we also run Algorithm 2 under the same noise level.
Figure 1: Interaction network.
The simulation results are illustrated in Figures 2 and 3. Figure 2 depicts the evolution of the system state. It can be seen that the variance of the system state trajectories under Algorithm 2 is much larger than those under Algorithm 4. This discrepancy arises from the direct integration of persistent noise into Algorithm 1, which sacrifices optimization accuracy and subsequently degrades the control performance of Algorithm 2. In our developed approach, the weakening factor $\chi^{k}$ is tailored to alleviate the influence of DP noise, ensuring accurate convergence of the system state. In addition, Figure 3 presents the evolution of the global constraint. It can be found that there exist constraint violations in Algorithm 2. However, owing to the implementation scheme developed in Section IV, our approach can guarantee the satisfaction of the global constraint.
Figure 2: System state evolution.
Figure 3: Global constraint evolution.
VI Conclusion
This paper developed a differentially private DMPC strategy for linear discrete-time systems with coupled global constraints. We showed that the DMPC method relying on the conventional distributed dual-gradient algorithm is susceptible to eavesdropping attacks. To address this issue, we incorporated a DP noise injection mechanism into the distributed dual-gradient algorithm, enabling privacy preservation while maintaining accurate optimization convergence. Furthermore, a practical implementation approach for DMPC was proposed, which guarantees the feasibility and stability of the closed-loop system. Simulation results validated the effectiveness of the developed privacy-preserving DMPC strategy. Future work will extend the differentially private framework for systems with uncertainties (e.g., robust and stochastic DMPCs).